# Danish Fungi 2024 DF20 Danish Fungi 2020

### v1

@@ Abstract
We introduce a new, highly challenging benchmark and a dataset – FungiTastic – based on data continuously collected over a twenty-year span. The dataset originates in fungal records labeled and curated by experts. It consists of about 350k multi-modal observations that include more than 650k photographs from 5k fine-grained categories and diverse accompanying information, e.g., acquisition metadata, satellite images, and body part segmentation. FungiTastic is the only benchmark that includes a test set with partially DNA-sequenced ground truth of unprecedented label reliability. The benchmark is designed to support (i) standard close-set classification, (ii) open-set classification, (iii) multi-modal classification, (iv) few-shot learning, (v) domain shift, and many more. We provide baseline methods tailored for almost all the use-cases. We provide a multitude of ready-to-use pre-trained models on HuggingFace and a framework for model training. A comprehensive documentation describing the dataset features and the baselines are available at GitHub and Kaggle.
@@ 1 Introduction
Biological problems provide a natural, challenging setting for benchmarking image classification methods. Consider the following aspects inherently present in biological data. The species distribution is typically seasonal and influenced by external factors such as recent precipitation levels. Species categorization is fine-grained, with high intra-class and inter-class variance. The distribution is often long-tailed; for rare species, only a very limited number of observations is available. New species are being discovered, raising the need for the “unknown” class option. Commonly, the set of classes has a hierarchical structure, and different misclassifications may have very different costs. Think of mistaking a poisonous mushroom for an edible one, which is potentially lethal, and an edible mushroom for a poisonous one, which at worse means coming back with an empty basket. Similarly, needlessly administering anti-venom after making a wrong decision about a harmless snake bite may be unpleasant, but its consequences are incomparable to not acting after a venomous bite.
@@
The properties of biological data listed above enable testing of, e.g., both open-set and closed-set categorization methods, robustness to prior and appearance domain shift, performance with limited training data, and dealing with non-standard losses. In contrast, most common benchmarks operate under the independent and identically distributed (i.i.d.) assumption, which is made valid by shuffling data and randomly splitting it for training and evaluation. In real-world applications, i.i.d data are rare since training data are collected well before deployment and everything changes over time [38].
@@
Data sources play an important role in benchmarking. In the age of LLMs and VLMs trained on possibly the entire content of the internet at a certain point in time, it is critical to have access to new, “unseen” data to guarantee that the tested methods are not evaluated on data they have indirectly “seen”, without knowing. Conveniently, many domains in nature are of interest to experts and the general public, who both provide a continuous stream of new and annotated data. The general public’s involvement introduces the problem of noisy training data; evaluating robustness to this phenomenon is also of practical importance.
In the paper, we introduce FungiTastic, a comprehensive multi-modal dataset of fungi observations which takes advantage of the favourable properties of natural data discussed above. The fungi observations include photographs, satellite images, meteorological observations, segmentation masks, and textual metadata. The metadata enrich the observations with attributes such as the timestamp, camera settings, GPS location, and information about the substrate, habitat, and biological taxonomy. By incorporating various modalities, the dataset support a robust benchmark for multi-modal classification, enabling the development and evaluation of sophisticated machine learning models under realistic and dynamic conditions.
@@
Classification of data originating in nature, including images of birds [3, 36], plants [14, 16], snakes [6, 25], and fungi [26, 35], has been used for benchmarking machine learning algorithms in several Fine-Grained Visual Categorization challenges; for a summary, see Table 1. Most of the commonly used datasets are small for current standards; the number of classes is also limited. The performance is often saturated, reaching total accuracy between 85–95 %; see the rightmost column of Tab. 1. Typically, the datasets are solely image-based and focused on traditional image classification; few of them offer basic attributes in metadata. Moreover, many popular datasets suffer from specific problems, e.g., regional, racial and gender biases [33], errors in labels [34, 4], and are saturated in accuracy.
@@
The key contributions of the proposed FungiTastic benchmark are:
@@
  • It includes diverse data types, such as photographs, satellite images, meteorological observations, segmentation masks, and textual metadata, providing a rich, multi-modal benchmark.
  • Each observation is annotated with attributes like timestamp, camera metadata, location (longitude, latitude, elevation), substrate, habitat, and biological taxonomy, facilitating detailed studies and advanced classification tasks.
@@
  • It addresses real-world challenges such as domain shifts, open-set, and few-shot classification, providing a realistic benchmark for developing robust machine learning models.
  • The dataset supports various evaluation protocols, including standard classification with novel-class detection, non-standard cost functions, time-sorted data for test-time adaptation methods, and few-shot classification.
  • The test and validation data have not been published before and thus remain unseen by large language models (LLMs) and vision-language models (VLMs), maintaining the integrity and robustness of the evaluation process.
user provided photographs (the knife left for scale)                  satellite image
Refer to caption Refer to caption Refer to caption Refer to caption   Refer to caption
Date: 2023-10-13 Habitat: Natural grassland Substrate: Soil Location: 56.84, 9.01 Taxon label: Agaricus fissuratus Elevation: 28.5m
Figure 1: An observation in FungiTastic includes one or more images of a specimen (three leftmost columns) and possibly some of its parts, such as the microscopic image of its spores (second from the right). Metadata available for virtually all observations are listed at the bottom. Geospatial information is available for all observations (right), DNA sequencing for a subset.
2 The FungiTastic Dataset
FungiTastic fungi is built on top of selected observations submitted to the Atlas of Danish Fungi before the end of 2023. Each observation includes at least one photograph and it is accompanied by additional metadata, see Figure1. In total, there are more than 650k images from 350k observations. The metadata include a multitude of attributes such as the timestamp, the camera settings, location (longitude, latitude, elevation), substrate, habitat, and taxonomy label. Not all observations have all of the attributes annotated, but the species attribute, which forms the basis for the primary classification task, has been annotated for all of the observations. Additionally, many images feature body-part segmentation masks are supplemented by satellite images or meteorological data.
Temporal division reflecting the natural seasonality in fungi distribution is provided to ensure a standardized approach for training and model evaluation. The FungiTastic–train dataset consists of all observations up to the end of 2021111the DF20 [26] training set with observations until the end of 2020 is a subset, the FungiTastic–val and FungiTastic–test datasets encompass all observations from 2022 and 2023, respectively.
We define two types of classes, "unknown," with no examples in the training set; the remaining classes are tagged "known". The unknown classes are used in evaluations of open-set recognition. The open-set classification tasks are challenging as many of the unknown species look similar to the known ones. The closed-set validation and test sets include only classes present in the training set.
Figure 2: Long-tailed distribution of classes (species) in the FungiTastic–M dataset sorted by frequency on the training set and color-coded by the set frequency, showing class prior shift between the training, test, and validation sets. The classes are sorted by their frequency on the training set. The number of classes in these sets are 215, 193, and 196, respectively. Best viewed in Zoom.
@@ 2 The FungiTastic Dataset
FungiTastic–M, where M is for mini, is a compact and challenging subset of the FungiTastic dataset consisting of all observations belonging to 6 hand-picked genera primarily targeted for prototyping. These genera form fruit bodies of the toadstool type with a large number of species. The FungiTastic–M comprises 46,842 images (25,786 observations) of 215 species, greatly reducing the computational requirements for training. The training, validation and test splits are the same as for the full dataset. The long-tailed class (species) distribution can be seen in Figure 2.
@@
FungiTastic–FS subset, FS for few-shot, is formed by species with less than 5 observations in the training set, which were removed from the main dataset. The subset contains 4,293 observations encompassing 7,819 images of a total of 2,427 species. As in the FungiTastic – closed set data, the split into validation and testing is done according to the year of acquisition.
Quantitative information about the FungiTastic is overviewed in Table 2.

### v2

@@ Abstract
We introduce a new, challenging benchmark and a dataset, FungiTastic, based on fungal records continuously collected over a twenty-year span. The dataset is labeled and curated by experts and consists of about 350k multimodal observations of 5k fine-grained categories (species). The fungi observations include photographs and additional data, e.g., meteorological and climatic data, satellite images, and body part segmentation masks. FungiTastic is one of the few benchmarks that include a test set with DNA-sequenced ground truth of unprecedented label reliability. The benchmark is designed to support (i) standard closed-set classification, (ii) open-set classification, (iii) multi-modal classification, (iv) few-shot learning, (v) domain shift, and many more. We provide tailored baselines for many use cases, a multitude of ready-to-use pre-trained models on HuggingFace, and a framework for model training. A comprehensive documentation describing the dataset features and the baselines are available at GitHub and Kaggle.
@@ 1 Introduction
Biological problems provide a natural, challenging setting for benchmarking image classification methods [40, 51, 52]. Consider the following aspects inherently present in biological data. The species distribution is typically seasonal and influenced by external factors such as recent precipitation levels. Species categorization is fine-grained, with high intra-class and inter-class variance. The distribution is often long-tailed; for rare species, only a few samples are available. New species are being discovered, raising the need for the “unknown” class option. Commonly, the set of classes has a hierarchical structure, and different misclassifications may have very different costs. Think of mistaking a poisonous mushroom for an edible one, potentially lethal, and an edible mushroom for a poisonous one, which at worst means returning home with an empty basket. Similarly, needlessly administering anti-venom after making a wrong decision about a harmless snake bite may be unpleasant, but its consequences are incomparable to not acting after a venomous bite.
@@
The properties of biological data listed above enable testing of, e.g., both open-set and closed-set categorization, robustness to prior and appearance domain shift, performance with limited training data, and dealing with non-standard losses. In contrast, most common benchmarks [10, 53] operate under the independent and identically distributed (i.i.d.) assumption, which is made valid by shuffling data and randomly splitting it for training and evaluation. In real-world applications, i.i.d data are rare since training data are collected well before deployment and everything changes over time [55].
@@
Data sources play an important role in benchmarking [19, 21]. In the age of LLMs and VLMs trained on possibly the entirety of the internet at a certain point in time, access to new, “unseen” data is crucial to ensure methods are tested on data not indirectly “seen”, without knowing. Conveniently, many domains in nature are of interest to experts and the general public, who both provide a continuous stream of new and annotated data [39, 49]. The general public’s involvement introduces the problem of noisy training data; evaluating robustness to this phenomenon is also of practical importance.
@@
Classification of data originating in nature, including images of birds [3, 53], plants [15, 18], snakes [6, 36], fungi [40, 52], and insects [17, 32] has widely been used for general benchmarking of machine learning algorithms, not just for fine-grained visual categorization. Most of the commonly used datasets are small for current standards; the number of classes is also limited. The performance is often saturated, reaching total accuracy between 85–95 %; see the rightmost column of Tab. 1. Typically, the datasets are solely image-based and focused on traditional image classification; few offer basic metadata attributes. Moreover, many popular datasets suffer from specific problems, e.g., regional, racial, and gender biases [48] and errors in labels [4, 50].
These limitations are particularly problematic for deployment, where machine learning models are expected to work with changing distributions and variable categories. As a result, the necessity of more comprehensive datasets that incorporate diverse modalities, reduce bias, and support more advanced learning paradigms, including multimodal, open-set, and few-shot classification, are becoming increasingly clear to better reflect the complexity of real-world data.
@@
In the paper, we introduce FungiTastic, a comprehensive multi-modal dataset of fungi observations [^1] A set of photographs and additional metadata describing one particular fungi specimen and surrounding environment. Usually, each photograph focuses on a different organ. For example observation, see Figure1 which takes advantage of the favorable properties of natural data discussed above. The fungi observations include photographs, satellite images, meteorological data, segmentation masks, and location-related metadata. The location metadata enriches the observations with attributes such as the timestamp, GPS location, and information about the substrate and habitat. By incorporating various modalities, the dataset supports a robust benchmark for multi-modal classification, enabling the development and evaluation of sophisticated machine-learning models under realistic and dynamic conditions. The key contributions of the proposed FungiTastic benchmark are:
@@
  • It includes diverse data types, such as photographs, satellite images, bioclimatic time-series data, segmentation masks, and contextual metadata (e.g., timestamp, camera metadata, location, substrate, and habitat), offering a rich, multimodal benchmark.
@@
  • It addresses real-world challenges such as domain shifts, open-set problems, and few-shot classification, providing a realistic benchmark for developing robust machine learning models.
  • The proposed benchmarks allow for addressing fundamental problems such as novel-class discovery, few-shot classification, and evaluation with non-standard cost functions.
User provided photographs Body part masks Satellite Climatic data
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Externally crawled metadata User provided observation metadata
 Toxicity: Edible Elevation: 28.5m Substrate: Soil Habitat: Nat. grassland
Bio zone: Atlantic Landcover: Grasslands Location: 56.84, 9.01 Date: 2023-10-13
Figure 1: A FungiTastic observation includes one or more photographs of an observed specimen and usually also a microscopic image of its spores. For a subset ($\sim$70k photographs), we provide body part segmentation masks. The user-provided observation metadata is available for virtually all observations as well as additional crawled metadata. Geospatial data is available for all observations, DNA sequencing for a subset. Last but not least, we provide climatic time-series data.
2 The FungiTastic Benchmark
FungiTastic is built from fungi observations submitted to the Atlas of Danish Fungi before the end of 2023, which were labeled by taxon experts on a species level. In total, more than 350k observations consisting of 650k photographs collected over 20 years are used. Apart from the photographs, each observation includes additional observation data (see Figure 1) ranging from satellite images, meteorological data, and tabular metadata (e.g., timestamp, GPS location, and information about the substrate and habitat) to segmentation masks and toxicity status. The vast majority of observations got most of the attributes annotated. For more details about the attribute description and its acquisition process, see Subsection 2.1.
Since the data comes from a long-term conservation project, its seasonality and naturally shifting distribution make it suitable for time-based splitting. In this so-called temporal division, all data collected before a certain point is used for training, while data collected after that point is reserved for evaluation. Specifically, in our case, the train datasets consist of all observations up to the end of 2021, and the val and test datasets include all observations from 2022 and 2023, respectively.
The FungiTastic benchmark is designed to support not just standard closed-set classification but also (i) open-set classification, (ii) few-shot learning, (iii) multi-modal classification, (iv) domain shift, and many more. For the most common task, we provide a short description of curated subsets below and additional quantitative information about all the subsets in the Appendix in Table 6.
FungiTastic is a general subset that includes most of the observations (around 340k) and species (i.e., 2,829) alongside all the additional observation data except for the body part masks. As a bonus, we include a DNA-based test set of 725 species and 2,041 observations. The FungiTastic comes along with dedicated validation and test sets specifically designed for closed-set and open-set scenarios. While the closed-set validation and test sets only include species present in the training set, the open-set also includes observations with species that were observed only after 2022 and 2023, i.e., species not available in the training set. All the species (i.e., categories) with no examples in the training set are labeled as "unknown".
@@ 2 The FungiTastic Benchmark
FungiTastic–Mini is a compact and challenging subset of the FungiTastic dataset designed primarily for prototyping and consisting of all observations belonging to 6 hand-picked genera (e.g., Russula, Boletus, Amanita, Clitocybe, Agaricus, and Mycena) [^2] These genera produce fruiting bodies of the toadstool type, which include many visually similar species and are of significant interest to humans due to their common use in gastronomy.. This subset comprises 46,842 images (25,786 observations) of 215 species, greatly reducing the computational requirements for training.
@@
FungiTastic–FS subset, FS for few-shot, is formed by species with less than 5 observations in the training set, which were removed from the main (FungiTastic) dataset. The subset contains 4,293 observations encompassing 7,819 images of a total of 2,427 species. As in the FungiTastic data, the split into validation and testing is done according to the year of acquisition.
2.1 Additional Observation Data
This section provides a comprehensive overview of the accompanying data available alongside the user-submitted photographs. For each subset, we describe the data itself and, if needed, its acquisition process as well. Below, we describe (i) tabular metadata, which include key environmental attributes and taxonomic information accompanying nearly all image observations, (ii) remote sensing data at fine-resolution geospatial scale for each observation site, (iii) meteorological data, which provides long-term climate variables, and (iv) body part segmentation masks that delineate specific morphological features of fungi fruiting bodies, such as caps, gills, pores, rings, and stems. All those metadata are integral to advancing research combining visual, textual, environmental, and taxonomic information.
Multi-band remote sensing data such as satellite images offer detailed and globally consistent environmental information at a fine resolution, making it a valuable resource for species categorization (i.e., identification) [44, 35]. To allow testing the potential of such data and to facilitate easy use of geospatial data, we provide multi-band (e.g., R, G, B, NIR, and elevation) satellite patches with 128$\times$128 pixel resolution (10m spatial resolution per pixel for R, G, B, and NIR, and 30m for elevation), centered on observation location. The data were extracted from rasters publicly available at Ecodatacube and ASTER Global Digital Elevation Model. While the extraction of the elevation data is straightforward, the raw pixel values in the Sentinel-2A rasters available on Ecodatacube might include extreme values. Therefore, we had to post-process raw data further to be in a standardized and expected form (i.e., an image with [0, 255] values). First, we clipped the values at 10,000. Next, the values were rescaled to a [0, 1] range and adjusted with a gamma correction factor of 2.5 (i.e., the values were raised to the power of 1/2.5). At last, the values were rounded and rescaled to [0, 255].
Body part segmentation masks of fungi fruiting body is essential for accurate identification and classification [9]. These morphological features provide crucial taxonomic information distinguishing some visually similar species. Therefore, we provide human-verified instance segmentation masks for all photographs in the FungiTastic–Mini (FungiTastic–M). We consider various semantic categories such as caps, gills, pores, rings, stems, etc. These annotations are expected to drive advances in interpretable recognition methods [43] and evaluation [22], with masks also enabling instance segmentation for separate foreground and background modeling [5]. All segmentation mask annotations were semi-automatically generated in CVAT using the Segment Anything Model [27].
Location-related metadata is provided for approximately 99.9% of the observations and describes the location, time, taxonomy and toxicity of the specimen, surrounding environment, and capturing device. See Table 2 for a detailed description of all available location-related metadata. While part of the metadata is usually provided by citizen scientists [^3] A member of the public who actively participates in scientific research and data collection, contributing valuable information and observations to support professional scientists. some attributes (e.g., elevation, land cover, and biogeographical) are crawled externally; all with a great potential to improve the classification accuracy and enable research on combining visual data with metadata.
Table 2: List of available location-related metadata. For a predominant number of observations, we provide a comprehensive set of data describing the surroundings or the specimen itself. Using such metadata for species identification allows to improve accuracy; see [11, 40].
Metadata Description
Habitat The environment where the specimen was observed. Selected from 32 values such as Mixed woodland, Deciduous woodland, etc.
Meteorological data and other climatic variables are vital assets for species identification and distribution modeling [1, 23]. In light of that, we provide 20 years of historical time-series monthly values of mean, minimum, and maximum temperature and total precipitation for all observations (see Figure 4 for example data). For each observation site, 20 years of data was extracted; for instance, an observation from the year 2000 includes data from 1980 to 2000. However, as the available climatic rasters only extend up to the year 2020, observations from 2020 to 2024 have missing values for those years not covered by existing data. In addition, we also provide 19 bioclimatic variables (related to temperature, seasonality, etc.) averaged over the period from 1981 to 2010. All data were extracted from climatic rasters available at CHELSA [25, 24].
Figure 4: Sample of available temperature data. First 5 years (out of 20 available years) of average, maximum, and minimum monthly temperatures for a randomly selected location.
3 FungiTastic Benchmarks
The diversity and unique features of the FungiTastic dataset allow for the evaluation of various fundamental computer vision and machine learning problems. We present several distinct benchmarks, each with its own evaluation protocol. This section provides a detailed description of each challenge and the corresponding evaluation metrics. Metrics are further defined in Appendix A.
Closed-set classification: The FungiTastic dataset is a challenging dataset with many visually similar species, heavy long-tailed distribution, and considerable distribution shifts over time. Since the fine-grained close-set classification methodology is well-defined, we follow the widely accepted standard, and we, apart from accuracy, use the macro-averaged F1-score ($\text{F}_{1}^{m}$).
Open-set classification: In the Atlas of Danish Fungi (our data source), new species are continuously added to the database, including previously unreported species. This ongoing data acquisition enables a yearly data split with a natural class distribution shift (see Figure 5), and many species in the test data are absent in the training set. We follow a widely accepted methodology, and we propose an AUC as the main metric. Besides, we calculate $\text{F}_{1}^{m}$ for two classes: "unknown" and "known."
Few-shot classification: All the categories with less than five samples in the FungiTastic, usually uncommon and rare species, form the few-shot subset. Being capable of recognizing those is of high interest to the expert. Since the few-shot dataset has no severe class imbalance like the other FungiTastic subsets, this benchmark’s main metric is top-1 accuracy. The F-1 score and top-k total accuracy are also reported. This challenge does not have any “unknown” category.
Chronological classification: Each observation in the FungiTastic (FungiTastic) dataset has a timestamp, allowing the study of species distribution changes over time. Fungi distribution is seasonal and influenced by weather, such as recent precipitation. New locations may be added over time, providing a real-world benchmark for domain adaptation methods, including online, continual, and test-time adaptation. The test dataset consists of fungi images ordered chronologically, meaning a model processing an observation at time $t$ can access all observations with timestamps $t^{\prime}<t$.
Classification beyond 0–1 loss function: Evaluation of classification networks is typically based on the 0-1 loss function, such as the mean accuracy, which applies to the metrics defined for the previous challenges as well. In practice, this often falls short of the desired metric since not all errors are equal. In this challenge, we define two practical scenarios: In the first scenario, confusing a poisonous species for an edible one (false positive edible mushroom) incurs a much higher cost than that of a false positive poisonous mushroom prediction. In the second scenario, the cost of not recognizing that an image belongs to a new species should be higher.
Segmentation: While acquiring human-annotated ground-truth segmentation masks can be resource-intensive, segmentation plays a critical role in advanced recognition methods and evaluation protocols for fine-grained image classification [5, 43]. High-accuracy segmentation of fungal images not only supports such methods but also enables automated analyses, such as assessing relationships between species-specific morphological features and environmental factors. This facilitates deeper insights into ecological and morphological patterns across locations, enhancing both model robustness and biological relevance. FungiTastic is suitable for evaluating semantic segmentation with the standard mean intersection over union (mIoU) metric and instance segmentation with the average precision (AP) metric.
Figure 5: Class distribution shift on the FungiTastic–M dataset.The long-term data acquisition captures a phenomenon related to natural changes in species presence, i.e., class prior shift. Sorted in descending order based on their occurrence in the training set (color-coded to indicate relative frequencies in the training, validation, and test sets). The training set includes data from before 2022 (215 species), the validation set from 2022 (196 species), and the test set from 2023 (193 species).
4 Baseline Experiments
In this section, we provide a variety of weak and strong baselines based on state-of-the-art architectures and methods for three FungiTastic benchmarks. Bellow, we report results only for the closed-set, few-shot learning, and zero-shot segmentation, but other baselines will be provided later in the supplementary materials, in the documentation, or on the dataset website.
4.1 Closed-set image classification
We train a variety of state-of-the-art CNN architectures to establish some baselines for closed-set classification on the FungiTastic and FungiTastic–M. All selected architectures were optimized with Stochastic Gradient Descent with momentum set to 0.9, SeeSaw loss [54], a mini-batch size of 64, and Random Augment [8] with a magnitude of 0.2. The initial learning rate was set to 0.01 (except for ResNet and ResNeXt, where we used LR=0.1), and it was scheduled based on validation loss.
Results: Similarly to other fine-grained benchmarks, while the number of params, complexity of the model, and training time are more or less the same, the transformer-based architectures achieved considerably better performance on both FungiTastic and FungiTastic–M and two different input sizes (see Table 3). The best-performing model, BEiT-Base/p16, achieved $\text{F}_{1}^{m}$ just around 40%, which shows the severe difficulty of the proposed benchmark.
Table 3: Closed-set fine-grained classification FungiTastic and FungiTastic–M. A set of selected state-of-the-art CNN- (top section) and Transformer-based (bottom section) architectures evaluated on the FungiTastic and FungiTastic–M test sets. All reported metrics show the challenging nature of the dataset. The best result for each metric is highlighted.
FungiTastic–M – 2242 FungiTastic – 2242 FungiTastic–M – 3842 FungiTastic – 3842
ResNet-50 61.7 79.3 35.2 62.4 77.3 32.8 66.3 82.9 39.8 66.9 80.9 36.3
ResNeXt-50 62.3 79.6 36.0 63.6 78.3 33.8 67.0 84.0 39.9 68.1 81.9 37.5
4.2 Few-shot image classification
Three baseline methods are implemented. The first baseline is standard classifier training with the Cross-Entropy (CE) loss. The other two baselines are nearest-neighbor classification and centroid prototype classification based on deep image embeddings extracted from large-scale pre trained vision models, namely CLIP [41], BioCLIP [47] and DINOv2 [33].
Standard deep classifiers are trained with the CE loss to output the class probabilities for each input sample. Nearest neighbors classification ($\mathbf{k}$-NN) constructs a database of training image embeddings. At test time, $k$ nearest neighbors are retrieved, and the classification decision is made based on the majority class of the nearest neighbors. Nearest-centroid-prototype classification constructs a prototype embedding for each class by aggregating the training data embeddings of the given class. The classification is performed based on the image embedding similarity to the class prototypes. These methods are inspired by prototype networks proposed in [45].
Results: While DINOv2 [33] embeddings greatly outperform CLIP [41] embeddings, BioCLIP [47] (CLIP finetuned on biological data) outperforms them both, highlighting the dominance of domain-specific models. Further, the centroid-prototype classification always outperforms the nearest-neighbor methods in terms of accuracy, while the nearest-neighbor wins over the centroid-prototype in F-score. Finally, the best standard classification models trained on the in-domain few-shot dataset underperform both DINOv2 and BioCLIP embeddings in the F-score, which shows the power of methods tailored to the few-shot setup. For results summary, refer to Table 4.
Table 4: Few shot classification on FungiTastic–Few-shot. (Left) – Pretrained deep descriptors with the nearest centroid and 1-NN classification. All pre-trained models are based on the ViT-B architecture, CLIP, and BioCLIP with patch size 32 and DINOv2 with patch size 16. (Right) – Standard classification with cross-entropy-loss. Evaluated on the FungiTastic–Few-shot test set.
[…]
CLIP 1-NN 6.1 2.8 4.2 5.3 9.9 11.5
centroid 7.2 2.2 13.0 3.6 7.4 12.5 12.5
DINOv2 1-NN 17.4 8.4 11.9 18.0 28.6 21.4
centroid 17.9 5.9 27.8 12.2 17.8 23.4 31.6
BioCLIP 1-NN 18.8 9.1 12.8 21.3 30.0 29.4
centroid 21.8 6.8 32.6 12.7 23.5 32.9 40.3
Architecture Input Top1   $\text{F}_{1}^{m}$ Top3
BEiT-B/p16 224$\times$224 11.0 2.1 17.4
384$\times$384 11.4 2.1 18.4
ConvNeXt-B 224$\times$224 14.0 2.7 23.1
384$\times$384 15.4 2.9 23.6
ViT-B/p16 224$\times$224 13.9 2.7 21.5
384$\times$384 19.5 3.7 29.0
4.3 Experiments with additional data
We provide baseline experiments using tabular metadata (habitat, month, substrate) based on our previous work [40]. The results in Table LABEL:tab:exps:tab_data show that all the attributes improve all the metrics. Individually, the addition of the habitat attribute results in the biggest gains (+2.27% of accuracy), followed by substrate (+1.15%) and month (+0.87%). Overall, habitat was the most efficient way to improve performance. With the combination of Habitat, Substrate, and Month, we improved the EfficientNet-B3 model’s performance on FungiTastic-Mini by 3.62%, 3.42% and 7.46% in Top1, Top3, and F1, respectively, indicating the gains are mostly orthogonal. Using the MetaSubstrate instead of Substrate resulted in performance lower by 0.2%, 0.5%, and 0.3% in Top1, Top3, and F1, respectively.
Table 5: Ablation on a combination of selected observation-related metadata. Utilizing a simple yet effective approach based on our previous work [40], we measure performance improvement using Habitat, Substrate, and Month and their combination. We also test how replacing Substrate with 31 variables for MetaSubstrate with nine variables with better interoperability to non-experts influences the performance. Evaluated with EfficientNet-B3 on FungiTastic-Mini test set.
Habitat
Month
Substrate
MetaSub.
Top1 69.5 +2.3 +0.9 +1.2 +0.9 +3.1 +3.0 +2.7 +1.9 +1.6 +3.6 +3.3 +3.5
$\text{F}_{1}^{m}$ 43.5 +4.0 +1.1 +2.3 +1.5 +6.0 +5.9 +5.1 +4.0 +3.2 +7.5 +6.8 +7.3
Top3 85.4 +2.3 +0.5 +0.8 +0.6 +2.7 +2.9 +2.6 +1.5 +1.1 +3.4 +3.1 +3.4
4.4 Segmentation
A zero-shot baseline for foreground-background binary segmentation of fungi is evaluated on the FungiTastic–M dataset. The method consists of two steps: 1. The GroundingDINO [30] (the ‘tiny’ version of the model) zero-shot object detection model is prompted with the text ‘mushroom’ and outputs a set of instance-level bounding boxes. 2. The bounding boxes from the first step are used as prompts for the SAM [27] segmentation model. All the experiments are conducted on images with the longest edge resized to 300 pixels while preserving the aspect ratio.
Results: The baseline method achieved an average per-image IoU of 89.36%. While the model exhibits strong zero-shot performance, it sometimes fails to detect mushrooms. These instances often involve small mushrooms, where a higher input resolution could enhance detection, and atypical mushrooms, such as very thin ones. Another common issue is SAM’s tendency to miss mushroom stems. The results for the simplified foreground-background segmentation task (rather than instance-level semantic segmentation) underline the need for further development of domain-specific models. Qualitative results, including random images and examples where the segmentation performs best and worst, are reported in Figures 8, 8, and 8.
Figure 6: Foreground-background segmentations on randomly selected FungiTastic–M photos. Highlighted pixels correspond to: true positives, false positives, and false negatives.
Figure 7: Foreground-background segmentations on FungiTastic–M with the highest IoU. Highlighted pixels correspond to: true positives.
Figure 8: Foreground-background segmentations on FungiTastic–M with the lowest IoU value. Highlighted pixels correspond to: false positives and false negatives.
5 Conclusion
In this work, we introduced the FungiTastic, a comprehensive and multi-modal dataset and benchmark. The dataset includes a variety of data types, such as photographs, satellite images, climatic data, segmentation masks, and observation metadata. FungiTastic has many interesting features, which make it attractive to the broad machine-learning community. With its data sampling spanning 20 years, precise labels, rich metadata, long-tailed distribution, distribution shifts over time, the visual similarity between the categories, and multimodal nature, it is a unique addition to the existing benchmarks.
In the provided baseline experiments, we demonstrate how challenging the FungiTastic Benchmarks are. Even state-of-the-art architectures and methods yield modest F-scores of 39.8% in closed-set classification and 9.1% in few-shot learning, highlighting the dataset’s difficulty compared to traditional benchmarks such as CUB-200-2011, Stanford Cars, and FGVC Aircraft. The proposed zero-shot baseline for the simplest segmentation task, binary segmentation of fungi fruiting body, achieved an average IoU of 89.36%, which still shows the potential for improvement in fine-grained visual segmentation of fungi.
Limitations mainly lie in the data collection process that influences data distribution. The majority of the data comes from Denmark, and further biases are potentially introduced by "sampling" in locations where certain species may be overrepresented due to their prevalence in frequently sampled areas or collector preferences. Additionally, not all meteorological data are available for every observation, which can affect of multi-modal classification approaches relying on such data.
Future work includes establishing continuous challenges to track progress across varied classification scenarios, expanding the set of weak and strong baselines, and including new test datasets for most benchmarks. Future efforts will also explore additional data, such as traits and textual descriptions of the species, to further improve the dataset’s multi-modal capabilities.
Online Resources
The complexity and size of the dataset require hosting the dataset in multiple spaces. Since the Kaggle dataset is allowing to store datasets just up to around 140GB we provide images for the vanilla FungiTastic dataset in smaller resolution. The full-size images, as well as the Croissant metadata, are available for download through dataset documentation.
Apart from the dataset, we also provide a wide variety of assets. Notably, we provide:
  • FGVC, a Pytorch-based framework designed for training deep learning classification models, hyperparameter fine-tuning, Weights&Biases logging, ONNX model export and HuggingFace deployment.
  • Sample notebooks that allows loading pre-trained models, prediction and making submission on Kaggle benchmarks.
  • All baseline pre-trained models available through the HuggingFace model hub allowing a straightforward integration and usability.
  • A comprehensive documentation with tables data samples and training strategies.
Acknowledgement
We extend our sincere gratitude to the mycologists from the Danish Mycological Society, particularly Jacob Heilmann-Clausen, Thomas Læssøe, Thomas Stjernegaard Jeppesen, Tobias Guldberg Frøslev, Ulrik Søchting, and Jens Henrik Petersen, for their invaluable contributions and expertise. We also thank the dedicated citizen scientists whose data and efforts have been instrumental to this project. Your support and collaboration have greatly enriched our work and made this research possible. Thank you for your commitment to advancing ecological understanding and conservation.
Appendix A Evaluation Metrics
A.1 Closed and open set classification
The main evaluation metric is $F$, the macro-averaged F1-score. For closed-set classification, the evaluation is standard, and for open-set, it is defined as
${F}_{K}=\frac{1}{{|K|}}\sum_{c\in{K}}F_{c},$ (2)
A.2 Classification beyond 0-1 loss function
For the classification beyond 0-1 cost, we follow the definition we set for the annual FungiCLEF competition [37, 38]. A metric of the following general form should be minimized.
$\mathcal{L}=\frac{1}{{N}}\sum_{i=1}^{{N}}{W}(y_{i},q_{1}(x_{i})),$ (4)
where $c_{u}=10$ and $c_{k}=1$.
Appendix B Supporting Figures and Tables